Can you help me break down the following equation? I have most of it but I am having a problem getting the fractions back into decimals once I have converted and solved the problem.
thank you, Melonie
-3/5[2/3 - 1/2(3/4-1/2)]

Answer by stanbon!(97)

Work with the parenthesis firs:
3/4-1/2= 3/4-2/4 = 1/4
Multiply this by (1/2) to get 1/8
Then 2/3-1/8 = (16-3)/24 = 13/24
Multiply by -3/5 to get -39/120
Cheers,
Stan H.